OpenStudy (questionaskerduh1243):

Check my work? Will medal and fan Researchers at jackson hospital asked 40 randomly selected patients about their sleep habits sixteen reported sleeping 8 hours or more each night if there are 700 patients in the hospital estimate the number patients who sleep 8 hours or more each night. 140 160 180 *** 280

OpenStudy (adrimit):

\[\frac{16}{40}=\frac{x}{700} \\x=4*70=280\]

OpenStudy (questionaskerduh1243):

Your correct

OpenStudy (questionaskerduh1243):

Thanks